OpenStudy (jakob):

Find the value of y when x equals 12. x+3y=21

OpenStudy (jakob):

@Torvela

OpenStudy (studygurl14):

Plug in 12 for x and solve for y

OpenStudy (jakob):

Well I now the first step

OpenStudy (studygurl14):

(12) + 3y = 21

OpenStudy (jakob):

no nevermind

OpenStudy (osprey):

9 y=(21-12)/3

OpenStudy (torvela):

ok the way i think of these is I almost do it in reverse so take 12+3y=21 then subtract 12

OpenStudy (torvela):

from both sides

OpenStudy (torvela):

then divide each side by 3

OpenStudy (torvela):

9 is wrong

OpenStudy (jakob):

-9

OpenStudy (torvela):

no

OpenStudy (jakob):

?

OpenStudy (torvela):

osprey had correct formula wrong answer

OpenStudy (osprey):

ok, sorry. my arithmetic's obviously bad today. I apologise for the inconvenience, and I wish this screen would stop jumping about ! take care

OpenStudy (torvela):

ok start by doing 21-12

OpenStudy (osprey):

21-12=9 DOESN'T IT ????

OpenStudy (jakob):

9

OpenStudy (torvela):

yes but then you have to do 9/3

OpenStudy (jakob):

so 3

OpenStudy (torvela):

yup

OpenStudy (jakob):

that my awnser?

OpenStudy (torvela):

now fill in y to check answer

OpenStudy (jakob):

Torvela help me with another?

OpenStudy (torvela):

so yes 3 is correct answer do you understand how i did it or should i re-explain?

OpenStudy (jakob):

I think i understand I am not sure.... I was that guy who didn want you helping that one question with TG sorry about that tho

OpenStudy (torvela):

it's ok

OpenStudy (osprey):

I'm just sea sick from the jumping about of the postings is all !

OpenStudy (torvela):

btw jakob i always thought about doing it backwards/opposite of what the equation is after filling in x

OpenStudy (torvela):

so since you know that when you fill in x it becomes 12+3y=21

OpenStudy (torvela):

next start with the normal number/ numbers without variables so in this case 12 and since its positive subtract it from 21

OpenStudy (torvela):

now you have 3y=9 and all you have to do is divide 9 by 3 and you have the answer